# Learning Where to Learn: Training Distribution Selection for Provable OOD Performance

### v1

@@ Appendix C Details for section 4Practical algorithms for training distribution design
Remark C.6 (overfitting and vanishing gradients).
A drawback of approximating the integral in the exact gradient (4.4) with a Monte Carlo average over the i.i.d. training samples in (C.6) and Line 4 of Alg. 3 is that the resulting gradient approximation incorporates the training error residual. If the trained model overfits to the data, then this residual becomes extremely small, which slows down the convergence of gradient descent significantly. In the extreme case, the gradient vanishes if the model perfectly fits the training data. Thus, the proposed method as currently formulated would likely fail for highly overparametrized models. To address this problem, one solution is to hold out a “validation” subset of the training data pairs for the adjoint state and Monte Carlo integral calculations; the remaining data should be used to train the model. Another approach is to normalize the size of the approximate gradient by its current magnitude. This adaptive normalization aims to avoid small gradients. In the numerical experiments of Sec. 5, we use kernel ridge regression with strictly positive regularization; this prevents interpolation of the data and helps avoid the vanishing gradient issue.
C.2 Derivations for subsection 4.2Upper-bound minimization
Recall the alternating minimization algorithm Alg. 2 from Sec. 4.2. We provide details about its numerical implementation. For gradient-based implementations, the first variation of the upper bound objective (4.8) is required. This is the content of Prop. 4.8, which we now prove.
Proof of Prop. 4.8.
Since the first term $\int_{\mathcal{U}}f(u)\nu(du)$ in the objective $F$ in (4.8) is linear in the argument $\nu$, the first variation of that term is simply $f$ itself [48, Def. 7.12, p. 262]. The chain rule of Euclidean calculus and a similar first variation calculation delivers the second term in the asserted derivative (4.9). The third and final term arising from the chain rule requires the first variation of $\nu\mapsto\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\mathsf{W}_{2}^{2}(\nu,\nu^{\prime})$. By [48, Prop. 7.17, p. 264, applied with cost $c(x,y)=\lVert x-y\rVert_{\mathcal{U}}^{2}/2$], this first variation equals the Kantorovich potential $2\phi_{\nu,\nu^{\prime}}$, where the asserted characterization of $\phi_{\nu,\nu^{\prime}}$ is due to Brenier’s theorem; see [11, Thm. 1.3.8, p. 27] or [54]. Finally, we exchange expectation over $\mathbb{Q}$ and differentiation using dominated convergence due to the assumed regularity of the set $\mathsf{P}$. ∎
The Wasserstein gradient of $F$ is useful for interacting particle discretizations. We have the following corollary of Prop. 4.8.
Corollary C.7 (Wasserstein gradient).
Instate the hypotheses of Prop. 4.8. The Wasserstein gradient $\nabla DF$ at $\nu\in\mathsf{P}\subseteq\mathscr{P}_{2}(\mathcal{U})$ is given by the function mapping any $u\in\mathcal{U}$ to the vector
@@
Lemma C.8 (chain rule: Gaussians).
Instate the hypotheses of Prop. 4.8. Suppose that $\mathcal{U}\subseteq\mathbb{R}^{d}$. Let $\nu_{\vartheta}\coloneqq\mathcal{N}(m_{\vartheta},\mathcal{C}_{\vartheta})$ and $\mathsf{F}(\vartheta)\coloneqq F(\nu_{\vartheta})$, where $F$ is given in (4.8). Write $p_{\vartheta}\colon\mathcal{U}\to\mathbb{R}_{\geq 0}$ for the density of $\nu_{\vartheta}$ with respect to any $\sigma$-finite dominating measure. Then for any parameter $\vartheta\in\mathcal{V}$, it holds that
@@
Furthermore, $\phi_{\nu,\nu^{\prime}}=\lVert{\,\cdot\,}\rVert_{\mathcal{U}}^{2}/2-\varphi_{\nu,\nu^{\prime}}$, where $\nabla\varphi_{\nu,\nu^{\prime}}=T_{\nu\to\nu^{\prime}}$ by Brenier’s theorem [11, Thm. 1.3.8, p. 27]. Upon integration, we find that
@@
for each $\vartheta$. Applying (4.9) from Prop. 4.8, closed-form formulas for the second moments of Gaussian distributions, and the Fubini–Tonelli theorem to exchange integrals completes the proof. ∎
Lem. C.8 requires the gradient $\nabla_{\vartheta}\log p_{\vartheta}$ of the log density with respect to the parameter $\vartheta$. In general, this gradient has no closed form and must be computed numerically, e.g., with automatic differentiation. However, when only the mean of the Gaussian family is allowed to vary, the following remark discusses an analytical expression for the gradient.
Remark C.9 (Gaussian mean parametrization).
Let $\mathcal{U}\subseteq\mathbb{R}^{d}$. If $\vartheta=m$ so that $\nu_{\vartheta}=\nu_{m}=\mathcal{N}(m,\mathcal{C}_{0})$ for fixed $\mathcal{C}_{0}$, then $\nabla_{m}\log p_{m}(u)=\mathcal{C}^{-1}(u-m)$, where $p_{m}$ is the Lebesgue density of $\nu_{m}$.
Appendix D Operator learning architectures
In this appendix, we provide detailed descriptions of the various deep operator learning architectures used in the paper. These are the DeepONet (SM D.1), NIO (SM D.2), and the newly proposed AMINO (SM D.3).
D.1 DeepONet: Deep Operator Network
@@
$\mathcal{G}_{\theta}(u)(z)=\big{\langle}\mathsf{B}\bigl{(}u(x^{(1)}),\ldots,u(x^{(J)})\bigr{)},\mathsf{T}(z)\big{\rangle}_{\mathbb{R}^{p}}=\sum_{k=1}^{p}\mathsf{B}_{k}\bigl{(}u(x^{(1)}),\ldots,u(x^{(J)})\bigr{)}\mathsf{T}_{k}(z)\,.$ (D.1)
During ERM training as in (2.1), the weights $\theta$ of both neural networks $\mathsf{B}$ and $\mathsf{T}$ are optimized together to minimize the average discrepancy between $\mathcal{G}_{\theta}(u)(z)$ and ground truth values $\mathcal{G}^{\star}(u)(z)$ over a dataset of input-output function pairs. This branch-trunk architecture enables DeepONet to approximate a broad class of operators, including the Neumann-to-Dirichlet (NtD) map in EIT and the conductivity-to-solution map in Darcy flow considered in this work; see Sec. 5 and SM E.
D.2 NIO: Neural Inverse Operator
Traditional approaches for constructing direct solvers for inverse problems, such as the D-bar method for EIT [50], have laid the foundation for direct solver approximations utilizing modern machine learning techniques. Among these, the Neural Inverse Operator (NIO) [39] framework stands out as a significant advancement. In the context of operator learning, many inverse problems can be formulated as a mapping from an operator $\Lambda_{a}$, dependent on a physical parameter $a$, to the parameter $a$ itself, i.e., $\Lambda_{a}\mapsto a$. In practical scenarios, the operator $\Lambda_{a}$ is not directly accessible; instead, we observe its action on a set of input-output function pairs $\{(f_{n},g_{n})\}_{n}$, where $g_{n}=\Lambda_{a}(f_{n})$. This observation leads to a two-step problem: first, infer the operator $\Lambda_{a}$ from the data pairs, and second, deduce the parameter $a$ from the inferred operator. This process can be further generalized by considering the operator as a pushforward map $(\Lambda_{a})_{\#}$, which transforms an input distribution $\mu$ of functions to an output distribution $\nu_{a}$ of functions, i.e., $(\Lambda_{a})_{\#}\mu=\nu_{a}$; see Sec. 2 for the definition of the pushforward operation. Consequently, a measure-centric machine learning approach to solving the inverse problem can be reformulated as learning the mapping $(\mu,\nu_{a})\mapsto a$.
@@ D.2 NIO: Neural Inverse Operator
D.3 AMINO: A Measure-theoretic Inverse Neural Operator
@@
Despite the innovative architecture of NIO, it exhibits certain limitations, particularly concerning its handling of input distributions. NIO primarily focuses on the output distribution $\nu_{a}$ during evaluation, neglecting variability and influence of the input distribution $\mu$. This oversight can lead to models that perform well on in-distribution (ID) data but lack robustness when faced with OOD scenarios. To address this shortcoming, we introduce A Measure-theoretic Inverse Neural Operator (AMINO), a framework that explicitly accounts for the input distribution in the inverse problem. By formulating the inverse task in a fully measure-theoretic setting and taking into account the input distribution, AMINO aims to improve OOD generalization. Like NIO, AMINO combines a DeepONet and an FNO. The key distinction is that AMINO incorporates both input and output functions as concatenated pairs $\{(f_{n},g_{n})\}_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}(\operatorname{Id}\times\Lambda_{a})_{\#}\mu$ which are passed to the branch network, as illustrated in Fig. 5.
@@
Figure 5: Schematic of the AMINO architecture, a variant of NIO, which combines a DeepONet, random sampling, and an FNO module. The model takes as input the grid point $z$ and paired function samples $\{(f_{n},g_{n})\}_{n}$ drawn from the joint distribution $(\operatorname{Id}\times\Lambda_{a})_{\#}\mu$ and outputs the target parameter $a$ at the point $z$.
NIO versus AMINO.
To compare the ID and OOD generalization error between NIO and AMINO, we start with the following forward problem setup taken from [39] for EIT on the domain $\Omega\coloneqq(0,1)^{2}\subset\mathbb{R}^{2}$. The model is the two-dimensional elliptic PDE boundary value problem
[…]
For one conductivity realization $a$ defined by
$x\mapsto a(x)=\exp\Biggl{(}\sum_{k=1}^{m}c_{k}\sin(k\pi x_{1})\sin(k\pi x_{2})\Biggr{)}\,,$ (D.3)
where $m=\lfloor\overline{m}\rfloor$, $\overline{m}\sim\mathsf{Unif}([1,5])$, and $\{c_{k}\}\sim\mathsf{Unif}([-1,1]^{m})$, we take $\mathsf{L}\in\mathbb{N}$ different Dirichlet boundary conditions $\{f_{\ell}\}_{\ell=1}^{\mathsf{L}}$ for $f$ in (D.2). These are given by
$x\mapsto f_{\ell}(x)=\cos\Bigl{(}\omega\bigl{(}x_{1}\cos(\theta_{\ell})+x_{2}\sin(\theta_{\ell})\bigr{)}\Bigr{)}$ (D.4)
with $\omega\sim\mathsf{Unif}([\pi,3\pi])$ and $\theta_{l}\coloneqq 2\pi\ell/\mathsf{L}$ for $\ell=1,\ldots,\mathsf{L}$. We denote by $\Lambda_{a}$ the DtN operator that maps the Dirichlet data of the PDE, $f_{\ell}$, to the Neumann boundary data, $a\frac{\partial u}{\partial\mathbf{n}}|_{\partial\Omega}$, where $u$ is the PDE solution and $\mathbf{n}$ is the outward normal unit vector. Here, the Dirichlet boundary conditions are functions drawn from a probability distribution $\mu$. AMINO learns the map from the joint Dirichlet–Neumann distribution $(\operatorname{Id}\times\Lambda_{a})_{\#}\mu$ to the conductivity $a$. In contrast, NIO only takes into account the associated Neumann data distribution $(\Lambda_{a})_{\#}\mu$ as its input, with the entire neural operator conditioned on a fixed Dirichlet data distribution $\mu$.
Figure 6: Training and testing performance of AMINO and NIO. (Top): Training error vs. iterations. (Middle) In-distribution testing error with $\omega\sim\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$, matching the training distribution performance. (Bottom) Out-of-distribution testing error with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$, differing from the training distribution. A rolling average with a window of $10$ epochs is applied to the middle and bottom plots. AMINO demonstrates improved out-of-distribution generalization performance compared to that of the NIO framework.
To compare ID generalization, both models are trained and tested on data sampled from the same Dirichlet distributions. The first two panels in Fig. 6 show that AMINO and NIO perform comparably in both training and ID settings. For OOD evaluation, the models are trained on Dirichlet data with $\omega\sim\mathsf{Unif}([5\pi,\frac{11\pi}{2}])$ and tested on data with $\omega\sim\mathsf{Unif}([\frac{\pi}{2},7\pi])$. The bottom panel in Fig. 6 demonstrates that AMINO achieves significantly better OOD generalization than NIO, reducing the OOD testing error by approximately half in this example.
Appendix E Details for section 5Numerical results
@@
This appendix describes detailed experimental setups for all numerical tests from Sec. 1 and Sec. 5 in the main text. Regarding function approximation, SM E.1 covers the bilevel gradient descent experiments. The remaining appendices concern operator learning. SM E.2 details the AMINO experiments. SM E.3 and SM E.4 describe the results of the alternating minimization algorithm when applied to learning linear NtD maps and the nonlinear solution operator of Darcy flow, respectively.
E.1 Bilevel minimization for function approximation
@@
We now describe the setting of the numerical results reported in Sec. 5 for Alg. 1 and its numerical version Alg. 3. Fix $m_{0}\in\mathcal{U}\coloneqq\mathbb{R}^{d}$. The goal of the experiments is to find the optimal Gaussian training distribution $\nu_{\vartheta}=\mathcal{N}(m_{0},\mathcal{C})\in\mathscr{P}_{2}(\mathbb{R}^{d})$ with respect to the bilevel OOD objective (4.1), specialized to the task of function approximation. We optimize over the covariance $\mathcal{C}$. For all $x\in\mathbb{R}^{d}$, the target test functions $\mathcal{G}^{\star}=g_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ for $i\in\{1,2,3,4\}$ that we consider are given by
$g_{1}(x)$ $=(1+0.1x_{3}^{4})\sin(x_{1})+7\sin^{2}(x_{2})\,,$ (E.1)
$g_{2}(x)$ $=\prod_{j=1}^{d}\frac{\lvert 4x_{j}-2\rvert+(j-2)/2}{1+(j-2)/2}\,,$ (E.2)
$g_{3}(x)$ $=10\sin(\pi x_{1}x_{2})+20(x_{3}-1/2)^{2}+10x_{4}+5x_{5}\,,\quad\text{and}\quad$ (E.3)
$g_{4}(x)$ $=\sqrt{(100x_{1})^{2}+\biggl{(}x_{3}(520\pi x_{2}+40\pi)-\frac{1}{(520\pi x_{2}+40\pi)(10x_{4}+1)}\biggr{)}^{2}}\,.$ (E.4)
@@
The first two functions come from [17] and the last two from [44, 47]. They are common benchmarks in the approximation theory literature; $g_{1}$ is known as the Ishigami function, $g_{2}$ the Sobol G function, $g_{3}$ the Friedmann-$1$ function, and $g_{4}$ the Friedmann-$2$ function. Although these functions are typically defined on compact domains (e.g., $[0,1]^{d}$ or $[-\pi,\pi]^{d}$), we instead consider approximating them on average with respect to Gaussian distributions (or mixtures thereof) on the whole of $\mathbb{R}^{d}$. We allow the dimension $d$ to be arbitrary; in particular, the functions $g_{1}$, $g_{3}$, and $g_{4}$ are constant in certain directions for large enough $d$.
We take $m_{0}=\frac{1}{2}(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$ in all setups except for the one concerning test function $g_{2}$, where we set $m_{0}=0$. Supplementary numerical results are given in Fig. 7, which reports the same information that Fig. 2 does, except now for the remaining functions $\{g_{i}\}$. The conclusions from these numerical results are the same. We remark that in Fig. 2, the leftmost subplot showing the training history is designed to highlight the long-time behavior (i.e., $1000$ steps) of the bilevel algorithm. In practice, we only run the algorithm for $O(10)$ to $O(100)$ gradient descent iterations; this is reflected in the leftmost column of Fig. 7, where $50$ steps are taken.
Figure 7: Bilevel Alg. 3 applied to the various ground truths $\{g_{i}\}$. (Left column) Evolution of $\mathsf{Err}$ from (E.8) over $50$ gradient descent iterations. (Center column) $\mathsf{Err}$ of model trained on $N$ samples from initial $\nu_{\vartheta}$ vs. optimal $\nu_{\vartheta}$. (Right column) Same as center, except incorporating the additional function evaluation cost incurred from Alg. 1. From top to bottom, each row represents target functions $\{g_{i}\}$ corresponding to those listed in each column of Tab. 1 (from left to right, excluding $g_{3}$ in the $d=5$ case). Shading represents two standard deviations away from the mean $\mathsf{Err}$ over $10$ independent runs.
@@
In all bilevel minimization experiments, we use the squared exponential kernel function $\kappa\colon\mathbb{R}^{d}\to\mathbb{R}^{d}\to\mathbb{R}_{>0}$ defined by
$(x,x^{\prime})\mapsto\kappa(x,x^{\prime})\coloneqq\exp\mathopen{}\mathclose{{}\left(-\frac{\lVert x-x^{\prime}\rVert^{2}_{2}}{\ell^{2}}}\right)\,,$ (E.5)
where $\ell>0$ is a scalar lengthscale hyperparameter. While not tuned for accuracy, we do adjust $\ell$ on the order of $O(1)$ to $O(10)$ depending on the target function $g_{i}$ to avoid blow up of the bilevel gradient descent iterations. We also experimented with the Laplace kernel $(x,x^{\prime})\mapsto\exp(-\lVert x-x^{\prime}\rVert_{1}/\ell)$; although giving similar accuracy as $\kappa$, the Laplace kernel led to slower bilevel optimization. Thus, we opted to report all results using the squared exponential kernel $\kappa$. Given a candidate set of training samples, the resulting kernel method is trained on these samples via kernel ridge regression as in Lem. C.4. We solve all linear systems with lower triangular direct solves using Cholesky factors of the regularized kernel matrices. No iterative optimization algorithms are required to train the kernel regressors. When fitting the kernel models to data outside of the bilevel gradient descent loop (see the center and right columns of Fig. 7), we set the ridge strength as $\sigma^{2}=10^{-3}/N$, where $N$ is the current training sample size.
@@
Moving on to describe the specific implementation details of Alg. 3, we initialize the covariance matrix as the identity $\vartheta^{(0)}=I_{d}$. We decrease the gradient descent step sizes $\{t_{k}\}$ in Alg. 3 according to a cosine annealing scheduler [35] with initial learning rate of $10^{-2}$ and final learning rate of 00. The regularization parameters $\{\sigma_{k}\}$ are also selected according to the cosine annealing schedule with initial value $10^{-3}$ and final value $10^{-7}$. Although the sequence $\{N_{k}\}$ of training sample sizes could also be scheduled, we opt to fix $N_{k}=250$ for all $k$ in Alg. 3 for simplicity (except for when approximating $g_{1}$, where we use $N_{k}=333$). No major differences were observed when choosing $\{N_{k}\}$ according to cosine annealing. The identification of more principled ways to select the sequences $\{\sigma_{k}\}$, $\{t_{k}\}$, and $\{N_{k}\}$ to optimally balance cost and accuracy is an important direction for future work.

### v3

@@ D.1 Necessity of finite data
Remark D.6 (overfitting and vanishing gradients).
A drawback of approximating the integral in the exact gradient (4.4) with a Monte Carlo average over the i.i.d. training samples in (D.11) and Line 4 of Alg. 3 is that the resulting gradient approximation incorporates the training error residual. If the trained model overfits to the data, then this residual becomes extremely small, which slows down the convergence of gradient descent significantly. In the extreme case, the gradient vanishes if the model perfectly fits the training data. Thus, the proposed method as currently formulated would likely fail for highly overparametrized models. To address this problem, one solution is to hold out a “validation” subset of the training data pairs for the adjoint state and Monte Carlo integral calculations; the remaining data should be used to train the model. Another approach is to normalize the size of the approximate gradient by its current magnitude. This adaptive normalization aims to avoid small gradients. In the numerical experiments of Sec. 5, we use kernel ridge regression with strictly positive regularization; this prevents interpolation of the data and helps avoid the vanishing gradient issue.
D.3 Derivations for subsection 4.2Upper-bound minimization
Recall the alternating minimization algorithm Alg. 2 from Sec. 4.2. We provide details about its numerical implementation. For gradient-based implementations, the first variation of the upper bound objective (4.8) is required. This is the content of Prop. 4.9, which we now prove.
Proof of Prop. 4.9.
Since the first term $\int_{\mathcal{U}}f(u)\nu(du)$ in the objective $F$ in (4.8) is linear in the argument $\nu$, the first variation of that term is simply $f$ itself [82, Def. 7.12, p. 262]. The chain rule of Euclidean calculus and a similar first variation calculation delivers the second term in the asserted derivative (4.9). The third and final term arising from the chain rule requires the first variation of $\nu\mapsto\operatorname{\mathbb{E}}_{\nu^{\prime}\sim\mathbb{Q}}\mathsf{W}_{2}^{2}(\nu,\nu^{\prime})$. By Santambrogio [82, Prop. 7.17, p. 264, applied with cost $c(x,y)=\lVert x-y\rVert_{\mathcal{U}}^{2}/2$], this first variation equals the Kantorovich potential $2\phi_{\nu,\nu^{\prime}}$, where the asserted characterization of $\phi_{\nu,\nu^{\prime}}$ is due to Brenier’s theorem; see Chewi [19, Thm. 1.3.8, p. 27] or [94]. Finally, we exchange expectation over $\mathbb{Q}$ and differentiation using dominated convergence due to the assumed regularity of the set $\mathsf{P}$. ∎
The Wasserstein gradient of $F$ is useful for interacting particle discretizations. We have the following corollary of Prop. 4.9.
Corollary D.7 (Wasserstein gradient).
Instate the hypotheses of Prop. 4.9. The Wasserstein gradient $\nabla DF$ at $\nu\in\mathsf{P}\subseteq\mathscr{P}_{2}(\mathcal{U})$ is given by the function mapping any $u\in\mathcal{U}$ to the vector
@@
Lemma D.8 (chain rule: Gaussians).
Instate the hypotheses of Prop. 4.9. Suppose that $\mathcal{U}\subseteq\mathbb{R}^{d}$. Let $\nu_{\vartheta}\coloneqq\mathcal{N}(m_{\vartheta},\mathcal{C}_{\vartheta})$ and $\mathsf{F}(\vartheta)\coloneqq F(\nu_{\vartheta})$, where $F$ is given in (4.8). Write $p_{\vartheta}\colon\mathcal{U}\to\mathbb{R}_{\geq 0}$ for the density of $\nu_{\vartheta}$ with respect to any $\sigma$-finite dominating measure. Then for any parameter $\vartheta\in\mathcal{V}$, it holds that
@@
Furthermore, $\phi_{\nu,\nu^{\prime}}=\lVert{\,\cdot\,}\rVert_{\mathcal{U}}^{2}/2-\varphi_{\nu,\nu^{\prime}}$, where $\nabla\varphi_{\nu,\nu^{\prime}}=T_{\nu\to\nu^{\prime}}$ by Brenier’s theorem [19, Thm. 1.3.8, p. 27]. Upon integration, we find that
@@
for each $\vartheta$. Applying (4.9) from Prop. 4.9, closed-form formulas for the second moments of Gaussian distributions, and the Fubini–Tonelli theorem to exchange integrals completes the proof. ∎
Lem. D.8 requires the gradient $\nabla_{\vartheta}\log p_{\vartheta}$ of the log density with respect to the parameter $\vartheta$. In general, this gradient has no closed form and must be computed numerically, e.g., with automatic differentiation. However, when only the mean of the Gaussian family is allowed to vary, the following remark discusses an analytical expression for the gradient.
Remark D.9 (Gaussian mean parametrization).
Let $\mathcal{U}\subseteq\mathbb{R}^{d}$. If $\vartheta=m$ so that $\nu_{\vartheta}=\nu_{m}=\mathcal{N}(m,\mathcal{C}_{0})$ for fixed $\mathcal{C}_{0}$, then $\nabla_{m}\log p_{m}(u)=\mathcal{C}_{0}^{-1}(u-m)$, where $p_{m}$ is the Lebesgue density of $\nu_{m}$.
Appendix E Operator learning architectures
In this appendix, we provide detailed descriptions of the various deep operator learning architectures used in the paper. These are the DeepONet (SM E.1), NIO (SM E.2), and the newly proposed AMINO (SM E.3).
E.1 DeepONet: Deep Operator Network
@@
$\mathcal{G}_{\theta}(u)(z)=\big\langle\mathsf{B}\bigl(u(x^{(1)}),\ldots,u(x^{(J)})\bigr),\mathsf{T}(z)\big\rangle_{\mathbb{R}^{p}}=\sum_{k=1}^{p}\mathsf{B}_{k}\bigl(u(x^{(1)}),\ldots,u(x^{(J)})\bigr)\mathsf{T}_{k}(z)\,.$ (E.1)
In practice, a constant bias vector is added to the preceding display. During ERM training as in (2.1), the weights $\theta$ of both neural networks $\mathsf{B}$ and $\mathsf{T}$ are optimized together to minimize the average discrepancy between $\mathcal{G}_{\theta}(u)(z)$ and ground truth values $\mathcal{G}^{\star}(u)(z)$ over a dataset of input-output function pairs. This branch-trunk architecture enables DeepONet to approximate a broad class of operators, including the Neumann-to-Dirichlet (NtD) map in EIT and the conductivity-to-solution map in Darcy flow considered in this work; see Sec. 5 and SM F.
E.2 NIO: Neural Inverse Operator
Traditional approaches for constructing direct solvers for inverse problems, such as the D-bar method for EIT [86], have laid the foundation for direct solver approximations utilizing modern machine learning techniques. Among these, the Neural Inverse Operator (NIO) [67] framework stands out as a significant advancement. In the context of operator learning, many inverse problems can be formulated as a mapping from an operator $\Lambda_{a}$, dependent on a physical parameter $a$, to the parameter $a$ itself, i.e., $\Lambda_{a}\mapsto a$. In practical scenarios, the operator $\Lambda_{a}$ is not directly accessible; instead, we observe its action on a set of input-output function pairs $\{(f_{n},g_{n})\}_{n}$, where $g_{n}=\Lambda_{a}(f_{n})$. This observation leads to a two-step problem: first, infer the operator $\Lambda_{a}$ from the data pairs, and second, deduce the parameter $a$ from the inferred operator. This process can be further generalized by considering the operator as a pushforward map $(\Lambda_{a})_{\#}$, which transforms an input distribution $\mu$ of functions to an output distribution $\nu_{a}$ of functions, i.e., $(\Lambda_{a})_{\#}\mu=\nu_{a}$; see Sec. 2 for the definition of the pushforward operation. Consequently, a measure-centric machine learning approach to solving the inverse problem can be reformulated as learning the mapping $(\mu,\nu_{a})\mapsto a$ [73, Sec. 2.4.1].
@@ E.2 NIO: Neural Inverse Operator
E.3 AMINO: A Measure-theoretic Inverse Neural Operator
@@
Despite the innovative architecture of NIO, it exhibits certain limitations, particularly concerning its handling of input distributions. NIO primarily focuses on the output distribution $\nu_{a}$ during evaluation, neglecting variability and influence of the input distribution $\mu$. This oversight can lead to models that perform well on in-distribution (ID) data but lack robustness when faced with OOD scenarios. To address this shortcoming, we introduce A Measure-theoretic Inverse Neural Operator (AMINO), a framework that explicitly accounts for the input distribution in the inverse problem. By formulating the inverse task in a fully measure-theoretic setting and taking into account the input distribution, AMINO aims to improve OOD accuracy. Like NIO, AMINO combines a DeepONet and an FNO. The key distinction is that AMINO incorporates both input and output functions as concatenated pairs $\{(f_{n},g_{n})\}_{n}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}(\operatorname{Id},\Lambda_{a})_{\#}\mu$ which are passed to the branch network, as illustrated in Fig. 7. See [73, Sec. 4.2.3] for a more detailed description of NIO and AMINO.
@@
Figure 7: Schematic of the AMINO architecture, a variant of NIO, which combines a DeepONet, averaging, and an FNO module. The model takes as input the paired function samples $\{(f_{\ell},g_{\ell})\}_{\ell=1}^{L}$ drawn from the joint distribution $(\operatorname{Id},\Lambda_{a})_{\#}\mu$ and outputs the target parameter $a$.
Appendix F Details for section 5Numerical results
@@
This appendix describes detailed experimental setups for all numerical tests from Sec. 1 and Sec. 5 in the main text. Regarding function approximation, SM F.1 covers the bilevel gradient descent experiments. The remaining appendices concern operator learning. SM F.2 details the AMINO experiments. SM F.3 and SM F.4 describe the results of the alternating minimization algorithm when applied to learning linear NtD maps and the nonlinear solution operator of Darcy flow, respectively. Last, SM F.5 presents results from applying nonparametric particle-based AMA to learn a parameter-to-solution map for the radiative transport equation.
F.1 Bilevel minimization for function approximation
@@
We now describe the setting of the numerical results reported in Sec. 5 for Alg. 1 and its numerical version Alg. 3. The goal of the experiments is to find the optimal Gaussian training distribution $\nu_{\vartheta}=\mathcal{N}(m,\mathcal{C})\in\mathscr{P}_{2}(\mathbb{R}^{d})$ with respect to the bilevel OOD objective (4.1), specialized to the task of function approximation. We optimize over the mean and covariance $\vartheta=(m,\mathcal{C})$. For all $x\in\mathbb{R}^{d}$, the target test functions $\mathcal{G}^{\star}=g_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ for $i\in\{1,2,3,4\}$ that we consider are given by
$g_{1}(x)$ $=\prod_{j=1}^{d}\frac{\lvert 4x_{j}-2\rvert+(j-2)/2}{1+(j-2)/2}\,$, (F.1a)
$g_{2}(x)$ $=10\sin(\pi x_{1}x_{2})+20(x_{3}-1/2)^{2}+10x_{4}+5x_{5}\,$, (F.1b)
$g_{3}(x)$ $=\sqrt{(100x_{1})^{2}+\biggl(x_{3}(520\pi x_{2}+40\pi)-\frac{1}{(520\pi x_{2}+40\pi)(10x_{4}+1)}\biggr)^{2}}\,,\quad\text{and}\quad$ (F.1c)
$g_{4}(x)$ $=\sum_{l=1}^{1000}c_{l}\kappa(x,x_{l})\,.$ (F.1d)
@@
The first function comes from [28] and the middle two from [77, 81]. They are common benchmarks in the approximation theory literature; $g_{1}$ is called the Sobol G function, $g_{2}$ the Friedmann-$1$ function, and $g_{3}$ the Friedmann-$2$ function. Although these functions are typically defined on compact domains (e.g., $0,1]^{d}$ or $\pi,\pi]^{d}$), we instead consider approximating them on average with respect to Gaussian distributions (or mixtures thereof) on the whole of $\mathbb{R}^{d}$. The final function $g_{4}$ is a linear combination of kernel sections for some kernel function $\kappa\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. We sample (and then fix) the coefficients $c_{l}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}([-1,1])$ and the centers $x_{l}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}([-4,4])^{\otimes d}$. We allow the dimension $d$ to be arbitrary for all test functions; in particular, the functions $g_{2}$ and and $g_{3}$ are constant in certain directions for large enough $d$.
@@
In all bilevel minimization experiments, we use the squared exponential kernel function $\kappa\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{>0}$ defined by
$(x,x^{\prime})\mapsto\kappa(x,x^{\prime})\coloneqq\exp\mathopen{}\mathclose{{\left(-\frac{\lVert x-x^{\prime}\rVert^{2}_{2}}{\ell^{2}}}}\right)\,$, (F.2)
where $\ell>0$ is a scalar lengthscale hyperparameter. While not tuned for accuracy, we do adjust $\ell$ on the order of $O(1)$ to $O(10)$ depending on the target function $g_{i}$ to avoid blow up of the bilevel gradient descent iterations. In particular, we set $\ell=1$ for $g_{1}$, $\ell=3$ for $g_{2}$, $\ell=2/1.1$ for $g_{3}$, and $\ell=5$ for $g_{4}$. We also experimented with the less regular Laplace kernel $(x,x^{\prime})\mapsto\exp(-\lVert x-x^{\prime}\rVert_{1}/\ell)$; although giving similar accuracy as $\kappa$, the Laplace kernel led to slower bilevel optimization. Thus, we opted to report all results using the squared exponential kernel $\kappa$. This same kernel is used to define $g_{4}$ in (F.1), which belongs to the RKHS of $\kappa$ as a result. Given a candidate set of training samples, the resulting kernel method is trained on these samples via kernel ridge regression as in Lem. D.4. We solve all linear systems with lower triangular direct solves using Cholesky factors of the regularized kernel matrices. No iterative optimization algorithms are required to train the kernel regressors. When fitting the kernel models to data outside of the bilevel gradient descent loop (see the center and right columns of Fig. 8, for instance), we set the ridge strength to be $\sigma^{2}=10^{-3}/N$, where $N$ is the current training sample size.
@@
Moving on to describe the specific implementation details of Alg. 3, we initialize $\vartheta^{(0)}=(m_{0},I_{d})$. We take $m_{0}=\frac{1}{2}(1,\ldots,1)^{\top}\in\mathbb{R}^{d}$ in all setups except for the one concerning test function $g_{1}$, where we set $m_{0}=(0,\ldots,0)^{\top}$. We decrease the gradient descent step sizes $\{t_{k}\}$ in Alg. 3 according to a cosine annealing scheduler [60] with initial learning rate of $10^{-2}$ and final learning rate of 0. The regularization parameters $\{\sigma_{k}\}$ are also selected according to the cosine annealing schedule with initial value $10^{-3}$ and final value $10^{-7}$. Although the sequence $\{N_{k}\}$ of training sample sizes could also be scheduled, we opt to fix $N_{k}=250$ for all $k$ in Alg. 3 for simplicity. No major differences were observed when choosing $\{N_{k}\}$ according to some cosine annealing schedule. The identification of more principled ways to select the sequences $\{\sigma_{k}\}$, $\{t_{k}\}$, and $\{N_{k}\}$ to optimally balance cost and accuracy is an important direction for future work.